A nutritionist wants to know the populatiin proportion of grade 11 students who eat vegetables. Pegged at a confidence of 95% survey among 1200 respondents was conducted and 200 said that they eat vegetables.




a.What is the margin of error?




b.Construct the confidence interval




c.Find the length of the confidence interval

1
Expert's answer
2022-06-16T10:01:13-0400

a.

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."


"ME=z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p}}{n}}"

"\hat{p}=\dfrac{x}{n}=\dfrac{200}{1200}=\dfrac{1}{6}"

"ME=1.96\times\sqrt{\dfrac{\dfrac{1}{6}(1-\dfrac{1}{6})}{1200}}\approx0.021"

b. The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-ME, \hat{p}+ME)"

"=(\dfrac{1}{6}-0.021, \dfrac{1}{6}+0.021)"

"=(0.146, 0.188)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.146 < p < 0.188," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.146, 0.188)."


c. The length of the confidence interval is "2ME=0.042."